What type of graph is:
6x^2-24x-5y^2-10y-11=0

Answer by richwmiller(17219)

hyperbola
hyperbola
foci | ((2-sqrt(11), -1) | (2+sqrt(11), -1))~~((-1.31662, -1) | (5.31662, -1))
vertices | ((2-sqrt(5), -1) | (2+sqrt(5), -1))~~((-0.236068, -1) | (4.23607, -1))
center | (2, -1)
semimajor axis length | sqrt(5)~~2.23607
semiminor axis length | sqrt(6)~~2.44949
focal parameter | 6/sqrt(11)~~1.80907
eccentricity | sqrt(11/5)~~1.48324
asymptotes | y = sqrt(6/5) x-2 sqrt(6/5)-1 | y = -sqrt(6/5) x+2 sqrt(6/5)-1
it can also be written
6(x-2)^2-5(y+1)^2-30 = 0